Chris is aged - y
Paul is twice the age of y
Peter is 3 years younger then Y
Their total age is 125, how old are they individually?

Answer by Cromlix(4381)

Hi there,
Chris is aged y
Paul is aged 2y
Peter is aged y - 3.
y + 2y + (y - 3) = 125
y + 2y + y - 3 = 125
Collect like terms
4y = 125 + 3
4y = 128
y = 32
Chris aged 32 years
Paul is aged 64 years
Peter is aged 29 years.
Hope this helps :-)